find the value of Z such that

i. the area of the right of Z is 0.2266

ii. the area between -Z and Z is 0.5722


1
Expert's answer
2021-06-22T09:12:03-0400

Solution:

(i) "P(Z>z)=0.2266"

"P(Z\le z)=1-P(Z>z)=1-0.2266=0.7734"

From z-score table:

"z=0.75"

(ii) "P(-z<Z<z)=0.5722"

"P(Z< z)-P(Z<-z)=P(Z<z)-[1-P(Z<z)]
\\=2P(Z<z)-1=0.5722"

"\Rightarrow 2P(Z<z)=1.5722
\\ \Rightarrow P(Z<z)=0.7861"

From z-score table:

"z=0.795"